OpenStudy (brooklyn):

an online movie rental company has 5 to the 3rd power action movie titles available to its members. the company also carries 5 to the 2nd power times as many drama movie titles. How many drama movie titles does the company carry?

OpenStudy (anonymous):

5^3 * 5^2 = 5 ^ (3+2 ) = 5 ^ 5 = 3125